# Enhancing selectivity using Wasserstein distance based reweighing

### v1

@@ 5.1 Bounding $1$-Wasserstein distance suffices
In this section, in Theorem 5.4, we show that the Wasserstein distance between two measures, corresponding in the sum of squares loss function, upper bounds the total variation distance between the invariant measures underlying the stochastic gradient descent (SGD) algorithms.
Assumption 5.1.
We assume that our input consists of $2n$ points among the vertices of the hypercube: $Q_{d(n)}:=\{0,1\}^{d(n)}$, where $d(n)$ is $\Theta(\log n)$.
Restricting the state-space to the hypercube is fairly standard in algorithms literature. For the drug discovery example, the state-space is just the binary molecule fingerprint vectors i.e., roughly $d(n)=2000$. Note that assuming training and test data-sets of equal size is more for clarity of presentation, as one can add dummy example points of 00 weight, if needed.
@@
Let $X\times Y$ denote the usual space of labeled examples i.e, in our case $X\subseteq\{0,1\}^{d(n)}$ is the set of feature values and $Y:=\{0,1\}$ is the set of labels. Our object of interest in this section is a neural network with smooth bounded activation functions. Let $y=f(w,x)$ denote the abstraction of our neural network, where $w$ denotes the real valued vector of weights. For a depth $p$ neural-net with polynomial activation functions of degree $q$, $f(w,x)$ is a polynomial in $x$ with degree at most $pq$.
@@
Recall that (see for example [6]), a stochastic gradient descent algorithm with loss function $\ell$ can be abstracted as the Itó diffusion in the limit of small step size:
@@
Assumption 5.2.
We assume that the diffusion corresponds to an uniformly elliptic generator, since that ensures the existence of a unique limiting (invariant) measure [2]. Furthermore, we assume $\sigma_{\mathcal{S}}$ is isotropic i.e, it’s a scalar multiple of the identity $\sigma\cdot\mathrm{Id}$ and that $\sigma_{\mathcal{S}}=\sigma_{\mathcal{T}}$, in Theorem 5.4.
We relax the isotropy assumption somewhat in a corollary (see supplement). Finally, for our situation of interest, i.e., $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})$ large, we make a covariate shift type assumption.
Assumption 5.3.
We assume that $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})=\Omega(1)$. Furthermore, $f$ and $y$ are bounded, say $y,f\in[0,1]$ and
Essentially, it says the data-sets have similar average loss in the same neighborhood for a given set of weights.
Theorem 5.4.
Suppose we train two neural networks, under the assumptions 5.2 and 5.3 above, on different input distributions, $\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}$, using the stochastic gradient descent (SGD) algorithm. Then, the total variation distance between their invariant measures can be bounded by $O(W_{1}(\mathbb{P}_{\mathcal{T}},\mathbb{P}_{\mathcal{S}}))$, in the limit as SGD step-size goes to 00.
Proof deferred to supplement.
Remark 5.5.
@@
5.2 The Greedy Algorithm
5.2.1 Reduction to bipartite matching
So far, we have established that $1$-Wasserstein metric is a sufficient topology to work with. This leads to the problem of computing the $1$-Wasserstein distance on two large datasets. That problem is equivalent to the minimum weight bipartite matching problem. In particular, we have the following lemma from [23].
Theorem 5.6.
[23] Given an instance of the optimal transport problem with supply and demands on two sets of points $(R,B)$, i.e, equivalently the $1$-Wasserstein distance computation problem in our case; we can construct an instance of the minimum weight bipartite matching problem such that solving the latter up to an approximation factor $\alpha$ will solve the former up to the same approximation factor $\alpha$.
The Algorithm ScaledGreedyWeight (below) carries out the reduction in Theorem 5.6 and calls GreedyMatch which matches two multisets embedded in a metric space using greedy algorithm on the edge lengths.
Notation: Scaling a discrete probability distribution $\mathbb{P}$ up by an integer factor of $C$ leads to a numerical rounding error of $\frac{1}{C\min\{\mathbb{P}\}}$, where ${}\min\{\mathbb{P}\}$ denotes the minimum positive value of density $\mathbb{P}$. Assume that we pick a large enough constant $C$ below, so that we can ignore the rounding error for the purposes of Theorem 5.13.
Algorithm 2 ScaledGreedyReweight (scale distributions and call bipartite matching)
@@ 5.2.1 Reduction to bipartite matching
2:Output: Probability distribution $\mathbb{P}_{B}^{\prime}$ supported on $B$. $\mathbb{P}_{B}^{\prime}$ is close to $\alpha\mathbb{P}_{R}+(1-\alpha)\mathbb{P}_{B}$ in $W_{1}$, under assumptions of Theorem 5.13.
3:for $r\in R$ do
@@
10:Use $\mathrm{GreedyMatch}(R^{\prime},B^{\prime})$ to compute the met (matched) demands, i.e., the extent to which the demands of $B$ that are actually fulfilled by $R$.
@@
5.2.2 Greedy algorithm and metric entropy
Recall that the data-points are set in the $d$-dimensional hypercube $Q_{d}$ with $\ell_{1}$ metric, where $d=\log^{O(1)}n$. The minimum weight bipartite matching problem is known to be harder than its non bipartite version. For example, the greedy algorithm is known to have a lower bound of $\Theta(n^{\log_{2}3/2})$ [20] for the bipartite version with $n$ data-set $\mathcal{T}$ and $n$ data-set $\mathcal{S}$ vertices. As an aside, a variant of the greedy algorithm (the hyper-greedy algorithm) provides a $\log n$ approximation in the non-bipartite case i.e, when any vertex can be matched to any other vertex. For the bipartite case, we obtain better approximation guarantees via the greedy algorithm, assuming small metric entropy of the input point sets.
Definition 5.7.
@@ 5.2.2 Greedy algorithm and metric entropy
In particular, Reingold and Tarjan [20] essentially show the following theorem.
Theorem 5.8.
[20] Given a set of $n$ data-set $\mathcal{T}$ and data-set $n$ $\mathcal{S}$ points in a metric space, the greedy algorithm returns a matching with weight that is within a factor of $\gamma^{\log_{2}\frac{3}{2}}$ of the minimum weight matching, where $\gamma$ is the length of the longest alternating cycle in the set (which can be $\Theta(n)$ for $Q_{\log n}$).
@@
Lemma 5.12 (Structural Lemma).
@@
Proof deferred to supplement. Note that an approximation factor of $d$ is trivial on $Q_{d}$ or on any set with $d_{\min}=1$ and $d_{\max}=d(n)$. The following corollary shows that the above indeed helps to improve upon the trivial bound for appropriately bounded instances.
Corollary 5.12.1.
Lemma 5.12 implies that the greedy algorithm achieves an approximation factor of $o(d^{3/4})$ on a $(d^{3/4},d^{3/4})$-bounded instance.
@@
We know that $N^{\mathrm{cov}}_{r}(E)\geq N^{\mathrm{ent}}_{r}(E)$ (see for example [9]). Therefore, Lemma 5.12 implies
@@
Theorem 5.13.
For $\eta=O(d^{\frac{1}{\xi\log_{2}3/2}})$, ($\xi>1$), Lemma 5.12 implies that the greedy algorithm achieves an approximation factor of $\max\{2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{\xi(1+\log_{2}(3/2))}}\right)\}$ on a $(\eta,\zeta)$-bounded minimum weight matching instance.
Together with Theorem 5.6, Theorem 5.13 implies that the greedy algorithm obtains the approximation factor on a $(\eta,\zeta)$-bounded Wasserstein distance computation instance. Proof deferred to supplement.
5.2.3 Small random samples suffice
In this subsection, we show that if the metric entropy is small, and so is the spread (see Definition 5.14) of the underlying distribution, then the empirical distribution of a much (polynomially) smaller sample is close to the actual distribution, in the 1-Wasserstein metric, with high probability.
Definition 5.14.
@@ 5.2.3 Small random samples suffice
${\scriptstyle\exists\ \alpha(n)\to 0,\ \lim_{n\to\infty}\frac{1}{\eta 2^{\zeta}}\ln\mathbb{P}\left(W_{1}(\hat{\mu}_{m},\mu)\geq\log\log n+o(1)S(\mu)\right)\leq-\Omega(1).}$ (7)
The proof is in the supplement. It closely follows the covering based proof of multidimensional Cramer’s theorem in its metric entropy version (exercise 6.2.19 in [9]) with two main differences: (1) we need the topology induced by the Wasserstein metric instead of the Levy metric, and (2) our space has large dimension, i.e., say $\log n$, which depends upon $n$ [^7] This is one reason why we can’t directly apply the result from [4]. The constants in the exponential in their theorems will depend on $n$, and it’s not immediately clear whether the dependence will lead to a non-trivial result.. The second point requires us to be more careful with the covering argument, and so we only prove a relatively weaker result, with the help of the transportation inequality from [5].
5.2.4 Efficiency guarantee: Greedy with random sampling
Theorems 5.155.13, and 5.6 imply the following efficiency guarantee about the greedy minimum weight bipartite matching algorithm (GreedyMatch) on a random sample, and therefore Algorithm 1 as well.
Theorem 5.16.
@@ 5.2.4 Efficiency guarantee: Greedy with random sampling
    $\mathcal{S}\cup\mathcal{T}$ admits a small covering: an $(\eta,\zeta)$ covering with $\eta,\zeta=O(\log^{c}n)$ and $\eta=O(\log^{c}n)$ for some constant $c\leq\frac{1}{1+\log_{2}(3/2)}$; and
@@
then the greedy algorithm achieves an approximation ratio of $\max\left(2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{1+\log_{2}(3/2)}}\right)\right)$ with probability $1-o(1)$, when computed on a small random sample of $r(n)$ fraction of data-points and $r(n)\to 0$.

### v2

@@ B.1 Choice of metric in Algorithm 1: Bounding $1$-Wasserstein distance suffices
In this subsection, in Theorem B.3, we show that the Wasserstein distance between two measures upper bounds the total variation distance between the invariant measures underlying the stochastic gradient descent (SGD) algorithms.
@@
Let $X\times Y$ denote the usual space of labeled examples i.e, in our case $X\subseteq\{0,1\}^{d(n)}$ is the set of feature values and $Y:=\{0,1\}$ is the set of labels. Our object of interest in this section is a neural network with smooth bounded activation functions. Let $y=f(w,x)$ denote the abstraction of our neural network, where $w$ denotes the real valued vector of weights. For a depth $p$ neural-net with piecewise polynomial activation functions of degree $q$, $f(w,x)$ is piecewise polynomial in $x$ with degree at most $pq$.
@@
We make a covariate shift type assumption.
Assumption B.1.
Assume that $f$ and $y$ are bounded, say $y,f\in[0,1]$, and for all $w$:
$\left|\mathbb{E}_{y\sim\mathbb{P}_{\mathcal{S}}(\cdot|x)}[(y-f(w,x))^{2}]-\mathbb{E}_{y\sim\mathbb{P}_{\mathcal{T}}(\cdot|x)}[(y-f(w,x))^{2}]\right|=O(1).$ (4)
Essentially, it says the data-sets have similar average loss in the same neighborhood for a given set of weights. It is also worth noting here that, if the invariant distribution of the SGD is concentrated around the local minima, i.e., if all but $\epsilon$ fraction of the mass of the invariant distribution, for some small positive $\epsilon$, are present within a small neighborhood around the local minima; then for the proof of Theorem 3.2 to remain valid, Assumption 3.1 only really needs to hold for $w$’s that are close to some locally optimal $w^{*}$.
Recall that, [^18] See for example Chaudhari and Soatto [2018] for an introductory discussion between SDEs and limiting SGD dynamics, and see Ben-Arous et al. [2023] for a more advanced discussion in this regard. a stochastic gradient descent algorithm with loss function $\ell$ can be abstracted as the Itó diffusion in the limit of small step size:
@@
Such a diffusion process is associated with an invariant measure or equilibrium distribution. In our context, this is the distribution of weight parameters of the neural net, as training time becomes very large. The existence of an invariant measure requires that the infinitesimal generator [^19] The infinitesimal generator $\mathcal{L}$ of the diffusion in Equation 5 may be written as: $\mathcal{L}f(w)=(\nabla_{w}f)^{T}(\nabla_{w}\ell_{w}(\mathbb{P}_{\mathcal{S}}))+\mathrm{Trace}\left((\nabla_{w}^{2}f(w))\frac{\sigma_{\mathcal{S}}\sigma_{\mathcal{S}}^{T}}{2}\right).$ It captures its most important properties, and in particular its adjoint $\mathcal{L}^{*}$ characterizes the invariant distribution (when it exists), i.e., $\mathcal{L}^{*}\rho=0$, where $\rho$ is the invariant distribution. associated with the diffusion be well behaved. In particular, Assumption B.2 about the generator ensures the existence of a unique limiting (invariant) measure, see Bianca and Dogbe [2017].
Assumption B.2.
We assume that the diffusion corresponds to an uniformly elliptic generator. Furthermore, we assume $\sigma_{\mathcal{S}}$ is isotropic i.e, it’s a scalar multiple of the identity $\sigma\cdot\mathrm{Id}$ and that $\sigma_{\mathcal{S}}=\sigma_{\mathcal{T}}$, in Theorem B.3.
We relax the isotropy assumption somewhat in Corollary C.1. Under the technical assumptions discussed above, we can prove the following result.
Theorem B.3.
Suppose we train two neural networks, under the assumptions B.2 and B.1, on different input distributions, $\mathbb{P}_{\mathcal{T}}$ and $\mathbb{P}_{\mathcal{S}}$, using the stochastic gradient descent (SGD) algorithm. If $W_{1}(\mathbb{P}_{\mathcal{S}},\mathbb{P}_{\mathcal{T}})=\Omega(1)$, then the total variation distance between their invariant measures can be bounded by $O(W_{1}(\mathbb{P}_{\mathcal{T}},\mathbb{P}_{\mathcal{S}}))$, [^20] Note that the $O(\cdot)$ here characterizes the linear dependence on $W_{1}(\mathbb{P}_{\mathcal{T}},\mathbb{P}_{\mathcal{S}})$, and there is a large constant factor hidden in the notation here due to the maximum eigenvalue of the inverse of the adjoint operator of the corresponding SDE. in the limit as training time goes to $\infty$ and SGD step-size goes to 00.
Proof deferred to Section C.
Remark B.4.
@@
B.2 The Greedy Algorithm
B.2.1 Reduction to bipartite matching
In this subsection, our main contribution is Theorem B.13. So far, we have established that $1$-Wasserstein metric is a sufficient topology to work with. This leads to the problem of computing the $1$-Wasserstein distance on two large datasets. That problem is equivalent to the minimum weight bipartite matching problem. In particular, we have the following lemma from Sharathkumar and Agarwal [2012].
Theorem B.5.
Sharathkumar and Agarwal [2012] Given an instance of the optimal transport problem with supply and demands on two sets of points $(R,B)$, i.e, equivalently the $1$-Wasserstein distance computation problem in our case; we can construct an instance of the minimum weight bipartite matching problem such that solving the latter up to an approximation factor $\alpha$ will solve the former up to the same approximation factor $\alpha$.
The GreedyAlgorithm (below) carries out the reduction in Theorem B.5 and calls GreedyMatch which matches two multisets embedded in a metric space using greedy algorithm on the edge lengths.
Notation: Scaling a discrete probability distribution $\mathbb{P}$ up by an integer factor of $C$ leads to a numerical rounding error of $\frac{1}{C\min\{\mathbb{P}\}}$, where ${}\min\{\mathbb{P}\}$ denotes the minimum positive value of density $\mathbb{P}$. Assume that we pick a large enough constant $C$ below, so that we can ignore the rounding error for the purposes of Theorem B.13.
Algorithm 2 GreedyAlgorithm
@@ B.2.1 Reduction to bipartite matching
2:  Output: Probability distribution $\mathbb{P}_{B}^{\prime}$ supported on $B$. Note $\mathbb{P}_{B}^{\prime}$ is close to $\alpha\mathbb{P}_{R}+(1-\alpha)\mathbb{P}_{B}$ in $W_{1}$, under assumptions of Theorem B.13$\triangleright$ Algorithm starts:
3:  For $r\in R$
@@
10:  Use the usual BFS (breadth first search) based greedy algorithm: GreedyMatch (Algorithm 3), on sets $R^{\prime}$ and $B^{\prime}$ to compute the met (matched) demands, i.e., the extent to which the demands of $B$ that are actually fulfilled by $R$.
@@
B.2.2 Greedy algorithm and metric entropy
Recall that the data-points are set in the $d$-dimensional hypercube $Q_{d}$ with $\ell_{1}$ metric, where $d=O(\log n)$. The minimum weight bipartite matching problem is known to be harder than its non bipartite version. For example, the greedy algorithm is known to have a lower bound of $\Theta(n^{\log_{2}3/2})$ Reingold and Tarjan [1981] for the bipartite version with $n$ data-set $\mathcal{T}$ and $n$ data-set $\mathcal{S}$ vertices.
Definition B.6.
@@ B.2.2 Greedy algorithm and metric entropy
In particular, Reingold and Tarjan Reingold and Tarjan [1981] essentially show the following theorem.
Theorem B.7.
Reingold and Tarjan [1981] Given a set of $n$ data-set $\mathcal{T}$ and data-set $n$ $\mathcal{S}$ points in a metric space, the greedy algorithm returns a matching with weight that is within a factor of $|\gamma|^{\log_{2}\frac{3}{2}}$ of the minimum weight matching, where $|\gamma|$ is the length of the longest alternating cycle $\gamma$ in the set (which can be $\Theta(n)$ for $Q_{\log n}$).
@@
Lemma B.11 (Structural Lemma).
@@
Proof deferred to Section C. Note that an approximation factor of $d$ is trivial on $Q_{d}$ or on any set with $d_{\min}=1$ and $d_{\max}=d(n)$. The following corollary shows that the above indeed helps to improve upon the trivial bound for appropriately bounded instances.
Corollary B.12.
Lemma B.11 implies that the greedy algorithm achieves an approximation factor of $o(d^{3/4})$ on a $(d^{3/4},d^{3/4})$-bounded instance.
@@
We know that $N^{\mathrm{cov}}_{r}(E)\geq N^{\mathrm{ent}}_{r}(E)$ (see for example Dembo and Zeitouni [2010]). Therefore, Lemma B.11 implies
@@
Theorem B.13.
For $\eta=O(d^{\frac{1}{\xi\log_{2}3/2}})$, ($\xi>1$), Lemma B.11 implies that the greedy algorithm achieves an approximation factor of $\max\{2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{\xi(1+\log_{2}(3/2))}}\right)\}$ on a $(\eta,\zeta)$-bounded minimum weight matching instance.
Together with Theorem B.5, Theorem B.13 implies that the greedy algorithm obtains the approximation factor on a $(\eta,\zeta)$-bounded Wasserstein distance computation instance. Proof deferred to Section C.
B.3 Small random samples suffice
In this subsection, our main contribution is Theorem B.15. we show that if the metric entropy is small, and so is the spread (see Definition B.14) of the underlying distribution, then the empirical distribution of a much (polynomially) smaller sample is close to the actual distribution, in the 1-Wasserstein metric, with high probability.
Definition B.14.
@@ B.3 Small random samples suffice
$\exists\ \alpha(n)\to 0,\ \lim_{n\to\infty}\frac{1}{\eta 2^{\zeta}}\ln\mathbb{P}\left(W_{1}(\hat{\mu}_{m},\mu)\geq\log\log n+o(1)S(\mu)\right)\leq-\Omega(1).$ (9)
The proof is in Section C. It closely follows the covering based proof of multidimensional Cramer’s theorem in its metric entropy version (exercise 6.2.19 in Dembo and Zeitouni [2010]) with two main differences: (1) we need the topology induced by the Wasserstein metric instead of the Levy metric, and (2) our space has large dimension, i.e., say $\log n$, which depends upon $n$. The second point requires us to be more careful with the covering argument, and so we only prove a relatively weaker result, with the help of the transportation inequality from Bolley and Villani [2005].
B.4 Greedy with random sampling
Theorems B.15B.13, and B.5 imply the following efficiency guarantee about the greedy minimum weight bipartite matching algorithm (GreedyMatch) on a random sample, and therefore Algorithm 1 as well.
Theorem B.16.
@@ B.4 Greedy with random sampling
    $\mathcal{S}\cup\mathcal{T}$ admits a small covering: an $(\eta,\zeta)$ covering with $\eta,\zeta=O(\log^{c}n)$ and $\eta=O(\log^{c}n)$ for some constant $c\leq\frac{1}{\xi(1+\log_{2}(3/2)})$, for any $\xi>1$; and
@@
then the greedy algorithm achieves an approximation ratio of $\max\left(2\zeta,O\left(d^{\frac{1+\xi\log_{2}(3/2)}{\xi(1+\log_{2}(3/2))}}\right)\right)$ with probability $1-o(1)$, when computed on a small random sample of $r(n)$ fraction of data-points and $r(n)\to 0$.